OpenStudy (anonymous):

Solve: square root of x+8=3

OpenStudy (lgbasallote):

sqrt (x+8) = 3? square first.. x+ 8 = 9 x = 9-8 x = 1

OpenStudy (anonymous):

No its this

OpenStudy (anonymous):

√(x+8)=3 Bring the square root to the right hand side and it becomes a square x + 8 = (3)2 x + 8 = 9 x = 9-8 x = 1